The following incomplete table shows a firm’s various costs of producing up to 6 units of output. Fill in as much of the table as possible. If you cannot determine the number in a box, explain why it is not possible to do so.


Q

TC

TVC

TFC

AC

MC

AVC

1

100

2

50

30

3

10

4

30

5

6

80
Answer #1
Q TC TVC TFC AC MC AVC
1 100 70 30 100 70
2 80 50 30 40 -20 25
3 90 60 30 30 10 20
4 150 120 30 37.5 60 30
5 30
6 30 80

TC=TFC+TVC

AC=TC/Q

MC=change in TC

AVC=VC/Q

It is not possible to compute the number in the other boxes as the information about TVC or MC is not given